A regression and correlation analysis resulted in the following information regarding a dependent variable (y) and an independent variable (x). Σx = 90 Σ(y - )(x - ) = 466 Σy = 170 Σ(x - )2 = 234 n = 10 Σ(y - )2 = 1434 ​ The F-statistic is Question 37 options: a18.341 b1.834 c14.673 dNone is correct.
Answer #1
SSx=Σx2-(Σx)2/n= 234
SSy=Σy2-(Σy)2/n= 1434
SP=Σxy-(ΣxΣy)/n= 466
SST=Syy= 1434
SSE =Syy-(Sxy)2/Sxx= 505.9829
SSR =(Sxy)2/Sxx = 928.0171
Source SS df MS F
regression 928.02 1 928.017 14.6727
error 505.98 8 63.248
total 1434.00 9

from above:F statistic =14.673